Independent t-Test: A researcher examines the short-term effect of a single treatment with acupuncture on cigarette...

  1. Independent t-Test: A researcher examines the short-term effect of a single treatment with acupuncture on cigarette smoking. Subjects who smoke ~ 1 ppd are assigned to one of two groups. One group (M2) receives acupuncture the second is a control group (M1) with no acupuncture. The number of cigarettes smoked the next day is recorded. You want to know if acupuncture will decrease cigarette smoking with alpha level of .05.

n= 8     M1 = 20.5        M2= 11.125     S21= 34.29       S22=57.84

  1. State you null and alternative hypothesis
  2. Find your critical region
  3. Calculate F-max
  4. Conclusion?
  5. Calculate t

Conclusion?
Answer #1

a) Null and alternative hypotheses

Ho:  1 = 2

H1:  1 > 2

b) For a = 0.05 and d.f = n1+ n2 -2

d.f = 8 + 8 -2 = 14

tcritical = t​​​​​​a , d.f = t​​​​​​0.05 , 14

tcritical = 1.761

c) F max = S​​​​​​2​​​​2​/S​​​​12

F max = 57.84/34.29

F max = 1.69

d) if t > 1.761 ,we reject the null hypothesis otherwise we fail to reject the null hypothesis.

e). t = M1 - M2/ Sp * sqrt(1/n1+1/n2)

Sp = sqrt((n1-1)s12 + (n2-1)s22)/n1+n2-2))

Sp = 6.79

t = 20.5 - 11.125/6.79*0.5

t = 2.76

Here t = 2.76 > 1.69 , so we reject the null hypothesis.

Conclusion : There is sufficient evidence to support the claim accupuncture treatment decreases cigarette smoking